C=100+0.8Yd
I=200
G=200
X=1000
M=20+0.2Y
(1) Equilibirium output
(2) Value of multiplier
1
Expert's answer
2020-04-03T09:48:59-0400


(1) let T = 100, than

Y=C+I+G+X-M

Y=100+0.8(Y-100)+200+200+1000-(20+0.2Y), Yd=Y-T

Y=100+0.8Y-80+200+200+1000-20-0.2Y

Y=1400+0.6Y

0.4Y=1400

Y=3500

(2) MPC=0.8

MPS=1-MPC=1-0.8=0.2

"M=\frac{1}{MPS}=\frac{1}{0.2}=5"